# An algorithmic version of the Hajnal–Szemerédi theorem

### v2

@@ Abstract.
A $K_{r}$-factor of a graph $G$ is a collection of vertex disjoint $r$-cliques covering $V(G)$. We prove the following algorithmic version of the classical Hajnal–Szemerédi Theorem in graph theory, when $r$ is considered as a constant. Given $r,c,n\in\mathbb{N}$ such that $n\in r\mathbb{N}$, let $G$ be an $n$-vertex graph with minimum degree at least $(1-1/r)n-c$. Then there is an algorithm with running time $2^{c^{O(1)}}n^{O(1)}$ that outputs either a $K_{r}$-factor of $G$ or a certificate showing that none exists, namely, this problem is fixed-parameter tractable in $c$. On the other hand, it is known that if $c=n^{\varepsilon}$ for fixed $\varepsilon\in(0,1)$, the problem is NP-C.
@@ 1. Introduction
In this paper we concern about the decision/search problem for the existence of certain subgraphs. We are interested in the target subgraphs whose order growing with the order of the host graph, as otherwise brute force search is enough. The most prominent example is the perfect matching problem, famously shown by Edmonds Blossom Algorithm [6] to be polynomial-time solvable. However, Hell and Kirkpatrick [15] showed that this is a rare example for spanning subgraphs – they showed that the decision problem for essentially all other spanning subgraphs in general graphs is NP-complete [^1] To be more precise, they showed that the $F$-factor problem is NP-complete if $F$ has a component on at least three vertices.. Therefore, it is natural to consider the restriction of the problem to certain class of host graphs so that the problem is tractable.
A classical line of research in extremal graph theory concerns about the smallest minimum degree condition or edge density condition that forces the existence of certain subgraphs. This has been a rich and active area of research since Dirac’s theorem on Hamiltonian cycles [5] proved in 1952. The decision problem restricts to this class of graphs (with large minimum degree) is trivial but the search problem is a nice and challenging problem.
Recently, Fomin, Golovach, Sagunov, and Simonov [8] started an interesting line of research at the intersection of complexity theory and graph theory. In particular, a classical theorem of Dirac [5] back to 1952 says that every $n$-vertex 2-connected graph $G$ with minimum degree $\delta(G)\geq 2$ contains a cycle of length at least $\min\{2\delta(G),n\}$. The authors of [8] proved an algorithmic extension of this result by showing that given $c\in\mathbb{N}$, there is an FPT algorithm parameterized in $c$ that determines whether such $G$ contains a cycle of length $\min\{2\delta(G)+c,n\}$, that is, the running time of the algorithm is $2^{O(c)}n^{O(1)}$  [^2] In general, a decision problem with parameter $c$ is in FPT parameterized in $c$ if it is solvable in time $f(c)\cdot n^{O(1)}$ for some computable function $f$.. A similar result on the Dirac’s theorem on Hamiltonicity is obtained earlier by Jansen, Kozma, and Nederlof [16]. In view of these results, a general question can be stated as follows.
@@
In this paper we take a further step towards Question 1 by studying an algorithmic extension of the celebrated Hajnal–Szemerédi Theorem. To state the theorem we first need some definitions. Given two graphs $F$ and $G$, an $F$-tiling in $G$ is a collection of vertex-disjoint copies of $F$ in $G$. The number of copies of $F$ is called the size of the $F$-tiling. An $F$-tiling is called perfect if it covers all the vertices of $G$. Perfect $F$-tilings are also referred to as $F$-factors.
@@
[13, Hajnal–Szemerédi] Every graph $G$ whose order $n$ is divisible by $r$ and whose minimum degree satisfies $\delta(G)\geq(1-1/r)n$ contains a $K_{r}$-factor.
@@
Following Question 1, an algorithmic extension of Theorem 1.1 could be the following.
Problem 1.2.
Given $r,c,n\in\mathbb{N}$ with $n\in r\mathbb{N}$ and a graph $G$ with $\delta(G)\geq(1-\tfrac{1}{r})n-c$, how difficult is it to decide whether $G$ has a $K_{r}$-factor?
Our main result resolves this problem when $r$ is considered as a constant [^3] Therefore, throughout this paper, when we use the big-O notation, the implicit constant depends only on $r$, not $c$ or $n$..
@@
Given $r,c,n\in\mathbb{N}$ with $n\in r\mathbb{N}$ and a graph $G$ with $\delta(G)\geq(1-\tfrac{1}{r})n-c$, there is an algorithm with running time $n^{O(1)}2^{O(c^{{r}^{r+4}})}$ that decides whether $G$ has a $K_{r}$-factor, that is, the problem is FPT parameterized in $c$. Moreover, the algorithm either returns a $K_{r}$-factor or a certificate showing that none exists.
By considering the complement graph, we have the equitable coloring version of Theorem 1.3.
Corollary 1.4.
Given $r,c,n\in\mathbb{N}$ with $n\in r\mathbb{N}$ and a graph $G$ with $\Delta(G)\leq\Delta+c$, there is an algorithm with running time $n^{O(1)}2^{O(c^{{r}^{r+4}})}$ that decides whether $G$ has an equitable $(\Delta+1)$-coloring, that is, the problem is FPT parameterized in $c$. Moreover, the algorithm either returns an equitable $(\Delta+1)$-coloring or a certificate showing that none exists.
@@
[21, Kühn–Osthus] Let $r,n\geq 3$ be integers and $\varepsilon\in(0,1)$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1-\frac{1}{r})n-n^{\varepsilon}$. Then the decision problem for $K_{r}$-factor in $G$ is NP-C.
Therefore, Theorem 1.3 bridges between this small gap and shows that the problem is FPT parameterized in $c$ (in particular, the problem is in P when $c=\text{poly}(\log n)$).
@@ 1.1. Related work
Fomin, Golovach, Sagunov, and Simonov [8] obtained a collection of results on existence of long cycles in 2-connected graphs with minimum degree assumptions or its slight weakening, including the algorithmic extension of Dirac’s Theorem mentioned earlier. In a followup paper, they also showed such an algorithmic extension of Erdős–Gallai Theorem, resolving an open problem. That is, they showed that for $c\in\mathbb{N}$, the decision problem for whether the circumference of a graph $G$ is at least $d(G)+c$ is FPT parameterized in $c$, where $d(G)$ is the average degree of $G$. See also [9, 10, 7] for further results.
Regarding Hamiltonicity, Dahlhaus, Hajnal, and Karpiński [4] showed that it is NP-C to decide whether an $n$-vertex graph $G$ with $\delta(G)\geq(1/2-\varepsilon)n$ contains a Hamilton cycle. For graphs $G$ with $\delta(G)\geq n/2-c$, Jansen, Kozma, and Nederlof [16] in 2019 gave an algorithm that runs in time $k^{c}n^{O(1)}$ for some constant $k$ which decides if $G$ contains a Hamilton cycle.
Another contribution to Question 1 is done by Han and Keevash [14], who studied Question 1 for minimum $(k-1)$-degree assumption forcing matchings of certain size in $k$-uniform hypergraphs. More precisely, they showed that for $c\in\mathbb{N}$, the decision problem for whether a $k$-uniform hypergraph $H$ containing a matching of size $\delta+c$ is FPT parameterized in $c$, where $\delta$ is the minimum $(k-1)$-degree of $H$.
@@ 1.2. Proof ideas
Let $G$ be an $n$-vertex graph $G$ with $\delta(G)\geq(1-1/r)n-c$. Then in time $n^{O(1)}2^{O(c^{{r}^{r+4}})}$ one can find a family $\mathcal{F}$ of $K_{r}$-tilings each of size $O(c^{r^{r+2}})$ defined on $V(G)$ such that $G$ has a $K_{r}$-factor if and only if there is $F\in\mathcal{F}$ such that $F\subseteq G$. Moreover, the existence of such $F$ in $G$ can be checked in time $n^{O(1)}2^{O(c^{{r}^{r+4}})}$, and the algorithm outputs a $K_{r}$-factor if one exists.
Indeed, we shall show that one can take $\mathcal{F}$ as a certain family of $K_{r}$-tilings of size $O(c^{r^{r+2}})$, which can be searched by the color-coding method. Then an immediate question will be how to make it a necessary condition. This is answered by our Lemma 1.7. Roughly speaking, it says that given such a rigid structure, one can conclude that if $G$ has a $K_{r}$-factor, then it has a small $K_{r}$-tiling of similar property.
@@
Let $r\geq r^{\prime}\geq 2$, $I\subseteq[r^{\prime}]$. Let $t_{1},t_{2},\ldots,t_{r^{\prime}}$ be $r^{\prime}$ integers such that $\sum_{i=1}^{r^{\prime}}t_{i}=0$ and $\sum_{i\in I}|t_{i}|>4r-3$. Let $G$ be an $n$-vertex graph and $(V_{1},V_{2},\dots,V_{r^{\prime}})$ be a partition of $V(G)$ with $|V_{i}|=n/r+t_{i}$ for each $i\in[r^{\prime}-1]$ and $|V_{r^{\prime}}|=(r-r^{\prime}+1)n/r+t_{r^{\prime}}$. Suppose that $G$ contains a $K_{r}$-factor $K$. Denote $t_{I}=\sum_{i\in I}|t_{i}|$. Then we can find a subset $K^{\prime}\subseteq K$ of size at most $(3t_{I})^{r}$ such that there exists $x\in\mathbb{N}$ with $|V_{i}\cap V(K^{\prime})|=x+t_{i}$, $i\in[r^{\prime}-1]\cap I$ and $|V_{r^{\prime}}\cap V(K^{\prime})|=(r-r^{\prime}+1)x+t_{r^{\prime}}$ if $r^{\prime}\in I$. In particular, $G-K^{\prime}$ still contains a $K_{r}$-factor.
@@
In our proof, we shall build such a vertex partition so that $t_{I}=t_{I}(c)=O(c^{r^{r+2}})$ and show that for almost all cases, this existence of $K^{\prime}$ is also a sufficient condition for the existence of $K_{r}$-factor; in some rare cases, a slightly more complicated condition works.
@@
For the remaining case ($G$ has at least one sparse $n/r$-set), it is crucial to know how many vertex-disjoint sparse $n/r$-sets $G$ contains. For this we use Algorithm 2 to find such a partition and edit it slightly. Let us denote the resulting vertex partition of $G$ by $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$ (see its definition in Section 6), where $s$ is the maximum number of vertex-disjoint sparse $n/r$-sets $G$ has. All the $A_{i}^{1}$ sets have size roughly $n/r$ (and thus $|B^{1}|\approx(r-s)n/r$). So let us call a copy of $K_{r}$ balanced if it contains exactly one vertex from each $A_{i}^{1}$ and $r-s$ vertices from $B^{1}$. Since each $A_{i}$ is sparse, we infer that each $G[A_{i},A_{j}]$ and $G[A_{i},B]$ is almost complete (and so do $G[A_{i}^{1},A_{j}^{1}]$ and $G[A_{i}^{1},B^{1}]$). This motivates the “ideal case” of our proof (Lemma 6.3). To reduce our $G$ to this ideal case, we need to (1) make the partition balanced, that is, its sizes have ratio $1:1:\cdots:1:(r-s)$ and (2) cover the vertices of wrong degree (in the sense of Lemma 6.3) by a small $K_{r}$-tiling with balanced copies. However, (1) is not always possible and indeed it reduces to finding a small $K_{r}$-tiling of certain type and we shall solve it by the color-coding technique developed by Alon–Yuster–Zwick. For a good summary on this strategy see our characterization results (Lemmas 6.11 and 6.21) in Section 6, which give the full version of Lemma 1.6.

### v3

# arXiv:2307.08056v3[math.CO] 27 Feb 2026

@@ Abstract.
A $K_{r}$-factor of a graph $G$ is a collection of vertex-disjoint $r$-cliques covering $V(G)$. We prove the following algorithmic version of the classical Hajnal–Szemerédi Theorem in graph theory, when $r$ is considered as a constant. Given $r,c,n\in\mathbb{N}$ such that $n\in r\mathbb{N}$, let $G$ be an $n$-vertex graph with minimum degree at least (11/r)nc(1-1/r)n-c. Then there is an algorithm with running time $2^{c^{O(1)}}n^{O(1)}$ that outputs either a $K_{r}$-factor of $G$ or a certificate showing that none exists, namely, this problem is fixed-parameter tractable in $c$. On the other hand, it is known that if $c=n^{\varepsilon}$ for fixed $\varepsilon\in(0,1)$, the problem is NP-C. By taking the complement, our result yields a similar result on the equitable $\Delta$-colorings of graphs of maximum degree $\Delta+c$, for $\Delta\in[n/r,n/(r-1)]$.
@@
L. Gan was partially supported by the National Natural Science Foundation of China (12401446) and the Fundamental Research Funds for the Central Universities
J. Han was partially supported by the National Natural Science Foundation of China (12371341)
J. Hu was partially supported by the National Natural Science Foundation of China (12301443)
@@ 1. Introduction
In this paper we concern about the decision/search problem for the existence of certain subgraphs. We are interested in the target subgraphs whose order grows with the order of the host graph, as otherwise brute force search is enough. The most prominent example is the perfect matching problem, famously shown by Edmonds Blossom Algorithm [5] to be polynomial-time solvable. However, Hell and Kirkpatrick [19] showed that this is a rare example for spanning subgraphs – they showed that the decision problem for essentially all other spanning subgraphs in general graphs is NP-C [^1] To be more precise, they showed that the $F$-factor problem is NP-C if $F$ has a component on at least three vertices.. Therefore, it is natural to consider the restriction of the problem to certain class of host graphs so that the problem is tractable.
A classical line of research in extremal graph theory concerns about the smallest minimum degree condition or edge density condition that forces the existence of certain subgraphs. This has been a rich and active research area since Dirac’s theorem on Hamiltonian cycles [4] proved in 1952. The decision problems restricted to this class of graphs (with large minimum degree) are trivial but the search problems for general subgraphs are nice and challenging problems.
A natural next step to these results is to study the decision problem in the graphs where the studied parameter is slightly off the threshold. In particular, restricted to the minimum degree case, this is when the minimum degree is strictly below the threshold. Indeed, this falls into the category of the above/below guarantee parameterization introduced by Mahajan, Raman and Sikdar [26], which was successfully applied in the study of several fundamental problems in parameterized complexity and kernelization. For illustrative examples and a comprehensive review, see the recent excellent survey by Gutin and Mnich [16].
In particular, a classical theorem of Dirac [4] back to 1952 says that every $n$-vertex 2-connected graph $G$ with minimum degree $\delta(G)\geq 2$ contains a cycle of length at least $\min\{2\delta(G),n\}$. Fomin, Golovach, Sagunov, and Simonov [10] in 2022 proved an algorithmic extension of this result by showing that given $c\in\mathbb{N}$, there is an FPT algorithm parameterized in $c$ that determines whether such $G$ contains a cycle of length $\min\{2\delta(G)+c,n\}$, that is, the running time of the algorithm is $^{O(c)}n^{O(1)} 22$2In general, a decision problem with parameter $c$ is in FPT parameterized in $c$ if it is solvable in time $f(c)\cdot n^{O(1)}$ for some computable function $f$.. A similar result on the Dirac’s theorem on Hamiltonicity is obtained earlier by Jansen, Kozma, and Nederlof [20] in 2019. In view of these results, a general question can be stated as follows.
@@
In this paper we take a further step towards Question 1 by studying an algorithmic extension of the celebrated Hajnal–Szemerédi Theorem. To state the theorem we first need some definitions. Given two graphs $F$ and $G$, an $F$-tiling (also called $F$-packing, $F$-matching in other contexts) in $G$ is a collection of vertex-disjoint copies of $F$ in $G$. The number of copies of $F$ is called the size of the $F$-tiling. An $F$-tiling is called perfect if it covers all the vertices of $G$. Perfect $F$-tilings are also referred to as $F$-factors.
@@
[17, Hajnal–Szemerédi] Every graph $G$ whose order $n$ is divisible by $r$ and whose minimum degree satisfies $\delta(G)\geq(1-1/r)n$ contains a $K_{r}$-factor.
@@
Following Question 1, the main goal of this paper is to study the following algorithmic problem stated in two equivalent forms.
Equitable $\Delta$-Coloring In Graphs Below Maximum DegreeParameter: $c$ Input: $n$-vertex graph $G$ with maximum degree $\Delta+c$. Question: Find an equitable $\Delta$-coloring in $G$ or report that it does not exist.
$K_{r}$-Factor In Graphs With High Minimum DegreeParameter: $c$ Input: $n$-vertex graph $G$ with minimum degree (11/r)nc(1-1/r)n-c. Question: Find a $K_{r}$-factor in $G$ or report that it does not exist.
Our main result deals with this problem when $r$ is considered as a constant [^3] Therefore, throughout this paper, when we use the big-$O$ notation, the implicit constant depends only on $r$, not $c$ or $n$..
@@
Given $r,c,n\in\mathbb{N}$ with $n\in r\mathbb{N}$ and an $n$-vertex graph $G$ with $\delta(G)\geq(1-\tfrac{1}{r})n-c$, there is an algorithm with running time $n^{O(1)}2^{O(c^{{r}^{r+4}})}$ that decides whether $G$ has a $K_{r}$-factor, that is, the problem is FPT parameterized in $c$. Moreover, the algorithm either returns a $K_{r}$-factor or a certificate showing that none exists. In particular, for $r=3$, the algorithm runs in time $n^{O(1)}2^{O(c)}$.
@@
[25, Kühn–Osthus] Let $r,n\geq 3$ be integers and $\varepsilon\in(0,1)$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1-\frac{1}{r})n-n^{\varepsilon}$. Then the decision problem for $K_{r}$-factor in $G$ is NP-C.
Therefore, Theorem 1.2 bridges between this small gap and shows that the problem is FPT parameterized in $c$ (in particular, the problem is in P when $c=O(\log^{r^{-(r+4)}}n)$).
A natural further extension is the algorithmic question on $K_{r}$-tilings in graphs of arbitrary size, which has been studied by Fellows, Knauer, Nishimura, Ragde, Rosamond, Stege, Thilikos, and Whitesides [8] who showed that (among others) in general graphs the search/decision problem for a $K_{r}$-tiling of size $c$ is FPT parameterized in $c$.
Theorem 1.4.
[8, Theorem 2] Given an $r$-vertex graph $F$ and $n$-vertex graph $G$ and integer $k>0$, there is an algorithm that determines whether $G$ has an $F$-tiling of size $k$ in time $O(n^{r}+2^{O(kr)})$.
It will be interesting to establish a result similar to Theorem 1.2 for $K_{r}$-tilings of arbitrary size. Together with a simple reduction to the factor case, our result yields the following result on the mixture of tilings.
Given a graph $G$, an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor in $G$ consists of a $K_{r}$-tiling of size $m$ and a $K_{r-1}$-tiling of size $m^{\prime}$ which are vertex-disjoint and cover all the vertices of $G$. Theorem 1.2 can be extended to decide whether a graph contains an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor.
Theorem 1.5.
Given $r,c,n,m,m^{\prime}\in\mathbb{N}$ such that $m\in[0,\frac{n}{r}]$, $m^{\prime}\in[0,\frac{n}{r-1}]$ and $mr+m^{\prime}(r-1)=n$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1-\frac{1}{r-1})n+\frac{m}{r-1}-c$. Then there is an algorithm with running time $n^{O(1)}2^{O(c^{r^{r+4}})}$ that decides whether $G$ contains an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor, that is, the problem is FPT parameterized in $c$. Moreover, the algorithm either returns an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor or a certificate showing that none exists.
Unfortunately, it does not solve the decision problem for $K_{r}$-tilings. More specifically, $G$ may have a $K_{r}$-tiling of size $m$ even if it has no $\{mK_{r},m^{\prime}K_{r-1}\}$-factor. Nevertheless, the complement version of Theorem 1.5 gives a nice consequence on equitable colorings (Corollary 1.6).
Proof of Theorem 1.5 using Theorem 1.2.
Let $A$ be a set of $m^{\prime}$ “omni” vertices disjoint with $V(G)$, and let $G^{\prime}$ be obtained by connecting all vertices of $A$ to all vertices of $G$. Note that the vertices of $A$ are independent. Note that $G^{\prime}$ contains a $K_{r}$-factor if and only if $G$ contains an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor. Indeed, if $G^{\prime}$ contains a $K_{r}$-factor $K$, then we remove the $m^{\prime}$ vertices of $A$ from $K$ and obtain an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor of $G$ (using the fact that the vertices in $A$ are independent, so no two of them are in the same $K_{r}$ in $K$). If $G$ contains an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor, then $G^{\prime}$ contains a $K_{r}$-factor because every vertex of $A$ connects to all vertices of $G$.
For every vertex $v\in V(G)$ we have
$d_{G^{\prime}}(v)\geq\delta(G)+m^{\prime}\geq(1-\frac{1}{r-1})n+\frac{m}{r-1}-c+m^{\prime}=(1-\frac{1}{r})(n+m^{\prime})-c$
and for every vertex of $A$ its degree is $n$ in $G^{\prime}$. Since $n\geq(1-\frac{1}{r})(n+m^{\prime})-c$, we have $\delta(G^{\prime})\geq(1-\frac{1}{r})(n+m^{\prime})-c.$ By applying Theorem 1.2 on $G^{\prime}$, the algorithm with running time $(n+m^{\prime})^{O(1)}2^{O(c^{r^{r+4}})}$ decides whether $G^{\prime}$ contains a $K_{r}$-factor. Hence the algorithm also decides whether $G$ contains an $\{mK_{r},m^{\prime}K_{r-1}\}$-factor. ∎
By considering the complement graph of $G$ and letting $\Delta:=m+m^{\prime}$ in Theorem 1.5, we have the following result on equitable colorings.
Corollary 1.6.
Given $r,c,n,\Delta\in\mathbb{N}$ such that $\frac{n}{r}\leq\Delta\leq\frac{n}{r-1}$. Let $G$ be an $n$-vertex graph with $\Delta(G)\leq\Delta+c$. Then there is an algorithm with running time $n^{O(1)}2^{O(c^{r^{r+4}})}$ that decides whether $G$ has an equitable $\Delta$-coloring, that is, the problem is FPT parameterized in $c$. Moreover, the algorithm either returns an equitable $\Delta$-coloring or a certificate showing that none exists.
@@ 1.1. Related work
Fomin, Golovach, Sagunov, and Simonov [10] obtained a collection of results on the existence of long cycles in 2-connected graphs with minimum degree assumptions or its slight weakening, including the algorithmic extension of Dirac’s Theorem mentioned earlier. In a followup paper, they [11] also showed such an algorithmic extension of Erdős–Gallai Theorem, resolving an open problem. That is, they showed that for $c\in\mathbb{N}$, the decision problem of whether the circumference of a graph $G$ is at least $d(G)+c$ is FPT parameterized in $c$, where $d(G)$ is the average degree of $G$. See also [9, 12, 13] for further results.
Regarding Hamiltonicity, Dahlhaus, Hajnal, and Karpiński [3] showed that it is NP-C to decide whether an $n$-vertex graph $G$ with $\delta(G)\geq(1/2-\varepsilon)n$ contains a Hamiltonian cycle. For graphs $G$ with $\delta(G)\geq n/2-c$, Jansen, Kozma, and Nederlof [20] in 2019 gave an algorithm that runs in time $2^{O(c)}n^{O(1)}$ which decides if $G$ contains a Hamiltonian cycle.
Another contribution to Question 1 is made by Han and Keevash [18], who studied Question 1 regarding the minimum (k1)(k-1)-degree assumption that forces matchings of certain size in $k$-uniform hypergraphs. More precisely, they showed that for $c\in\mathbb{N}$, the decision problem of whether a $k$-uniform hypergraph $H$ containing a matching of size $\delta+c$ is FPT parameterized in $c$, where $\delta$ is the minimum (k1)(k-1)-degree of $H$.
@@ 1.2. Proof ideas
Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1-1/r)n-c$. Then in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ (resp. $n^{O(1)}2^{O(c)}$ for $r=3$) one can find a family $\mathcal{F}$ of $K_{r}$-tilings each of size $O(c^{r^{r+2}})$ (resp. $O(c)$ for $r=3$) defined on $V(G)$ such that $G$ has a $K_{r}$-factor if and only if there is $F\in\mathcal{F}$ such that $F\subseteq G$. Moreover, the existence of such $F$ in $G$ can be checked in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ (resp. $n^{O(1)}2^{O(c)}$ for $r=3$), and the algorithm outputs a $K_{r}$-factor if one exists.
Indeed, we shall show that one can take $\mathcal{F}$ as a certain family of $K_{r}$-tilings of size $O(c^{r^{r+2}})$ (resp. $O(c)$ for $r=3$), which can be searched by the color-coding method. Then an immediate question will be how to make it a necessary condition. This is answered by our Lemma 1.8. Roughly speaking, it says that given such a rigid structure, one can conclude that if $G$ has a $K_{r}$-factor, then it has a small $K_{r}$-tiling of similar property.
@@
Let $r\geq r^{\prime}\geq 2$, $I\subseteq[r^{\prime}]$. Let $t_{1},t_{2},\ldots,t_{r^{\prime}}$ be $r^{\prime}$ integers such that $\sum_{i=1}^{r^{\prime}}t_{i}=0$ and $\sum_{i\in I}|t_{i}|>4r-3$. Let $G$ be an $n$-vertex graph and $(V_{1},V_{2},\dots,V_{r^{\prime}})$ be a partition of $V(G)$ with $|V_{i}|=n/r+t_{i}$ for each $i\in[r^{\prime}-1]$ and $|V_{r^{\prime}}|=(r-r^{\prime}+1)n/r+t_{r^{\prime}}$. Suppose that $G$ contains a $K_{r}$-factor $K$. Denote $t_{I}=\sum_{i\in I}|t_{i}|$. Then we can find a subset $K^{\prime}\subseteq K$ of size at most $g(t_{I})$ such that there exists $x\in\mathbb{N}$ with $|V_{i}\cap V(K^{\prime})|=x+t_{i}$, $i\in[r^{\prime}-1]\cap I$ and $|V_{r^{\prime}}\cap V(K^{\prime})|=(r-r^{\prime}+1)x+t_{r^{\prime}}$ if $r^{\prime}\in I$, where $g(t_{I})=(3t_{I})^{r}$ for $r\geq 4$ and $g(t_{I})=8t_{I}$ for $r=3$. In particular, $G-K^{\prime}$ still contains a $K_{r}$-factor.
@@
In our proof, we shall build such a vertex partition so that $t_{I}=t_{I}(c)=O(c^{r^{r+2}})$ (resp. $O(c)$ for $r=3$) and show that for almost all cases, this existence of $K^{\prime}$ is also a sufficient condition for the existence of $K_{r}$-factor; in some rare cases, a slightly more complicated condition works.
@@
For the remaining case ($G$ has at least one sparse $n/r$-set), it is crucial to know how many vertex-disjoint sparse $n/r$-sets $G$ contains. For this we use Algorithm 2 to find such a partition $(A_{1},\dots,A_{s},B)$ and edit it slightly. Let us denote the resulting vertex partition of $G$ by $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$ (see its definition in Section 6), where $s$ is the maximum number of vertex-disjoint sparse $n/r$-sets $G$ has. All the $A_{i}^{1}$ sets have size roughly $n/r$ (and thus $|B^{1}|\approx(r-s)n/r$). So let us call a copy of $K_{r}$ balanced if it contains exactly one vertex from each $A_{i}^{1}$ and rsr-s vertices from $B^{1}$. Since each $A_{i}$ is sparse, we infer that each $G[A_{i},A_{j}]$ and $G[A_{i},B]$ is almost complete (and so do $G[A_{i}^{1},A_{j}^{1}]$ and $G[A_{i}^{1},B^{1}]$). This motivates the “ideal case” of our proof (Lemma 6.3). To reduce our $G$ to this ideal case, we need to (1) make the partition balanced, that is, its sizes have ratio $1:1:\cdots:1:(r-s)$ and (2) cover the vertices with wrong degree (in the sense of Lemma 6.3) by a small $K_{r}$-tiling with balanced copies. However, (1) is not always possible and indeed it reduces to finding a small $K_{r}$-tiling of certain type and we shall solve it by the color-coding technique developed by Alon–Yuster–Zwick. For a good summary on this strategy see our characterization results (Lemmas 6.11 and 6.21) in Section 6, which give the full version of Lemma 1.7.